# Optimization of the Downlink Spectral- and Energy-Efficiency of RIS-aided Multi-user URLLC MIMO Systems

### v1

@@ II-B Signal Model
${\bf x}_{l}=\sum_{k=1}^{K}{\bf x}_{lk}\in\mathbb{C}^{N_{BS}\times 1},$ (10)
[…]
@@
$=\!\!\underbrace{{\bf H}_{lk,l}{\bf x}_{lk}}_{\text{Desired Signal}}\!+\!\!\underbrace{\sum_{j\neq k}{\bf H}_{lk,l}{\bf x}_{lj}}_{\text{Intra-cell Interference}}\!+\!\underbrace{\sum_{i\neq l}{\bf H}_{lk,i}{\bf x}_{i}}_{\text{Intercell Interference}}\!+\!\underbrace{{\bf n}_{lk}}_{\text{Noise}},$
[…]
@@ II-C Channel Dispersion, Rate and EE Expressions
A MIMO channel can be modeled as a set of parallel AWGN channels, and [24, Theorem 78] can be employed to obtain the achievable rate of MIMO channels associated with FBL coding. Note that [24, Theorem 78] is based on the optimal power allocation for a point-to-point MIMO communication link; however, the FBL rate expressions can be formulated for any arbitrary power allocation as shown in [24, Section 4.5.4]. In the following lemma, we calculate the achievable FBL rate of users, when the interference is treated as noise for decoding the corresponding signal at the receivers.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
which quantifies how energy efficient the network is.
@@ II-D Problem Statement
where constraint (23c) can be interpreted as a latency constraint for each user, as discussed in [11, Sec. II.D]. Moreover, functions $f_{i}$, $\forall i$ are linear functions of the rates/EEs and/or transmit/receive powers. The general problem in (II-D) can include an extensive range of practical optimization scenarios including the maximization of the minimum weighted rate, sum rate, global EE and minimum EE. We refer the reader to [43, Sec. II.B] for more discussions on the format of the functions $f_{i}$s as well as of the family of optimization problems that can be formulated as (II-D).
@@ III Proposed optimization framework
In this section, we propose iterative schemes for solving (II-D) by leveraging AO, MM-based, and FP algorithms. Specifically, we first fix the RIS coefficients to $\{{\bf\Psi}^{(t-1)}\}$ and update the transmit covariance matrices as $\{{\bf P}^{(t)}\}$ by solving (II-D). We then alternate and update the RIS coefficients, while $\{{\bf P}\}$ is fixed to $\{{\bf P}^{(t)}\}$. We iterate this procedure until convergence is reached. Unfortunately, the optimization problems are non-convex and complicated even when the RIS elements (or covariance matrices) are fixed. Thus, we propose a suboptimal scheme based on MM to solve the corresponding problems. Below, we present our solutions for updating the transmit covariance matrices and RIS elements in separate subsections.
@@ III-A Updating Transmit Covariance Matrices
To update $\{{\bf P}\}$, we introduce a new set of variables $\{{\bf Q}\}=\{{\bf Q}_{lk}\}_{\forall lk}$, where ${\bf Q}_{lk}$ is a positive semi-definite matrix and ${\bf P}_{lk}={\bf Q}_{lk}{\bf Q}_{lk}^{H}$. Equivalently, we can compute ${\bf Q}_{lk}$ as ${\bf P}_{lk}^{1/2}$. To attain a suboptimal solution for (II-D), we leverage an MM-based technique. More specifically, we first obtain suitable concave surrogate functions for the rates. Then, we update ${\bf P}_{lk}^{(t)}={\bf Q}_{lk}^{(t)}{\bf Q}_{lk}^{(t)^{H}}$, $\forall l,k$, by solving the corresponding surrogate optimization problems. To derive concave lower bounds for the FBL rates, we utilize the bounds in the following lemmas.
@@
$r_{lk}\geq\tilde{r}_{lk}=a_{lk}+2\sum_{ij}\mathfrak{R}\left\{\text{{\em Tr}}\left({\bf A}_{lk,ij}\mathbf{Q}_{ij}^{H}\bar{\mathbf{H}}_{lk,i}^{H}\right)\right\}\\ -\text{{\em Tr}}\left({\bf B}_{lk}(\bar{\bf H}_{lk,l}\mathbf{Q}_{lk}\mathbf{Q}_{lk}^{H}\bar{\bf H}_{lk,l}^{H}+\mathbf{D}_{lk})\right)$ (27)
@@
$a_{lk}$ $\!=\!\ln\left|\!{\bf I}\!+\!\bar{\bf D}_{lk}^{-1}\bar{\bf S}_{lk}\!\right|\!\!-\!\!\text{{\em Tr}}\!\left(\bar{\bf D}_{lk}^{-1}\bar{\bf S}_{lk}\right)\!\!-\!\frac{Q^{-1}\!(\epsilon)\!(\!\sqrt{\bar{V}_{lk}}\!\!+\!2I)}{2\sqrt{n_{t}}},$
@@
${\bf B}_{lk}$ $\!=\bar{\mathbf{D}}^{-1}_{lk}\!-\!(\bar{\mathbf{S}}_{lk}+\bar{\mathbf{D}}_{lk})^{-1}\!\!+\!\frac{Q^{-1}(\epsilon)}{\sqrt{n_{t}\bar{V}_{lk}}}\bar{\mathbf{D}}_{lk}(\bar{\mathbf{S}}_{lk}+\bar{\mathbf{D}}_{lk})^{-2}\!,$
@@
[…]
@@
which is a fractional function of $\{{\bf Q}\}$ with a concave numerator and convex denominator. Note that although the surrogate lower bounds for the rates in Lemma 5 are concave, the $\tilde{f}_{i}$s are not necessarily concave, since they might be a linear function of the EE metrics. Substituting the $f_{i}$s by the $\tilde{f}_{i}$s leads to
@@
The optimization problem (III-A) is convex for the maximization of the minimum and/or sum rates. Hence, it can be efficiently solved by existing numerical tools. However, (III-A) is non-convex for GEE maximization as well as for the maximization of the minimum weighted EE of the users, since the EE and/or GEE functions are not concave in $\{{\bf Q}\}$. Fortunately, a globally optimal solution of the minimum weighted EE of the users and/or GEE maximization problems can be obtained by Dinkelbach-based algorithms, since the numerator of $\tilde{e}_{lk}$ (or $\tilde{g}$) is concave, while its denominator is convex. In the following, we provide a globally optimal solution of (III-A) for the maximization of the minimum weighted EE of the users and the maximization of the GEE.
@@ III-A1 Maximization of the Minimum EE
In this case, (III-A) can be written as
@@
Upon employing the generalized Dinkelbach algorithm (GDA), we can derive the globally optimal solution of (III-A1) by iteratively solving the convex optimization problem [42]
@@ III-A2 Maximization of the GEE
In this case, (III-A) is equivalent to
@@
III-A3 Discussion on Single-stream Data Transmission
In this case, the rank of the matrix ${\bf P}_{lk}$ for all $l,k$ is equal to one. It means that the matrix ${\bf P}_{lk}$ can be written as ${\bf P}_{lk}={\bf q}_{lk}{\bf q}_{lk}^{H}$, where ${\bf q}_{lk}\in\mathbb{C}^{N_{BS}\times 1}$. In other words, when single-stream data transmission is employed, the BSs perform only a beamforming to transmit data, and we have to optimize the beamforming vectors instead of transmit covariance matrices. To this end, we only have to replace the matrices ${\bf Q}_{lk}$ by the vectors ${\bf q}_{lk}$ and employ the lower bounds in Lemma 5. Indeed, the schemes proposed in this subsection can be applied to both single- and multiple-stream data transmission schemes. Note that as indicated before, the maximum number of streams per users is $\min(N_{BS},N_{u})$. Thus, if the transmitter and/or the receiver have a single antenna, we are restricted to single-stream data transmissions.
@@ III-B Optimizing the RIS Elements
Now, we update $\{\bf{\Psi}\}$ by solving (II-D) for fixed $\{{\bf P}^{(t)}\}$, i.e.,
@@
which is non-convex since the rates are not concave in $\{\bf{\Psi}\}$ and the set $\mathcal{T}$ might be non-convex. In the following, we first consider regular RISs and then, describe how the solution can be applied to the more sophisticated STAR-RIS, utilizing the MS scheme. To find a suboptimal solution for (III-B), we leverage an approach based on MM. Specifically, we first obtain a suitable concave lower bound for the rates and then, convexify $\mathcal{T}$ if it is not already a convex set. Since the rates have similar structures in $\{{\bf Q}\}$ and in $\{\bf{\Psi}\}$, we can utilize the concave lower bounds in Lemma 5 to attain concave lower bounds for the rates and construct suitable surrogate optimization problems for updating $\{\bf{\Psi}\}$. In the subsequent corollary, we provide the concave lower bounds for the rates.The proof of this corollary closely resembles that of Lemma 5, and thus, we omit it here.

### v2

@@ II-B Signal Model
${\bf x}_{l}=\sum_{k=1}^{K_{l}}{\bf x}_{lk}\in\mathbb{C}^{N_{BS,l}\times 1},$ (10)
where ${\bf x}_{lk}$ is the signal intended for user Ulk, which is a zero-mean complex Gaussian random vector with covariance ${\bf P}_{lk}=\mathbb{E}\{{\bf x}_{lk}{\bf x}_{lk}^{H}\}$, where $\mathbb{E}\{{\bf x}\}$ is the mathematical expectation of ${\bf x}$. We assume that the zero-mean signals ${\bf x}_{lk}$ are independent from each other, i.e., $\mathbb{E}\{{\bf x}_{lk}{\bf x}_{ij}^{H}\}={\bf 0}$ for $i\neq l$ and/or $j\neq k$. Additionally, we denote the covariance matrix of ${\bf x}_{l}$ by ${\bf P}_{l}=\mathbb{E}\{{\bf x}_{l}{\bf x}_{l}^{H}\}$. Since the signals ${\bf x}_{lk}$ are zero-mean and independent random vectors for all $l$ and $k$, we have $\mathbf{P}_{l}=\sum_{k}\mathbf{P}_{lk}$. The set containing all the feasible transmit covariance matrices is denoted as $\mathcal{P}$ and it is given by
@@
$=\!\!\underbrace{{\bf H}_{lk,l}{\bf x}_{lk}}_{\text{Desired Signal}}\!+\!\!\underbrace{\sum_{j\neq k}{\bf H}_{lk,l}{\bf x}_{lj}}_{\text{Intracell Interference}}\!+\!\underbrace{\sum_{i\neq l}{\bf H}_{lk,i}{\bf x}_{i}}_{\text{Intercell Interference}}\!+\!\underbrace{{\bf n}_{lk}}_{\text{Noise}},$ (12)
where ${\bf n}_{lk}$ is the zero-mean additive white Gaussian noise at Ulk with covariance matrix $\sigma^{2}{\bf I}$, where ${\bf I}$ denotes the identity matrix. Note that, in (12), the differences between the intercell and intracell links are carefully taken into account, and each user is indeed affected by both intercell and intracell interference. In this paper, we treat interference as noise, which is optimal for maximizing the sum rate [50] or the generalized degree of freedom [51] when the interference is weak. An alternative strategy for treating interference as noise (TIN) is to detect and cancel interference, which is known as successive interference cancellation (SIC), and it is optimal when the interference is strong. SIC requires more advanced user devices as well as more sophisticated signaling design. Moreover, to detect and cancel interference at the users takes some time, which may lead to violating the latency constraint in URLLC-related applications.
Note that we employ Gaussian signaling in this work similar to most studies of wireless communication systems both with and without RISs. In practice, typically discrete constellations are employed. Studies based on Gaussian signaling are nevertheless important since they provide valuable insights into the system performance and represent an upper bound for the performance of the technologies studied. Additionally, there are studies on the comparison of Gaussian signals and discrete constellations, e.g., [52, 53, 54]. The performance gap between the discrete constellations and Gaussian signals grows as the number of bits/symbols in the discrete constellations increases, but it eventually saturates. To account for this performance gap and the idealized assumption, one can employ a signal-to-noise ratio (SNR) offset.
@@ II-C Channel Dispersion, Rate and EE Expressions
A MIMO channel can be modeled as a set of parallel AWGN channels, and [27, Theorem 78] can be employed to obtain the achievable rate of MIMO channels associated with FBL coding. Note that [27, Theorem 78] is based on the optimal power allocation for a point-to-point MIMO communication link; however, the FBL rate expressions can be formulated for any arbitrary power allocation as shown in [27, Section 4.5.4]. In the following lemma, we calculate the achievable FBL rate of users, when the interference is treated as noise for decoding the corresponding signal at the receivers.
@@
The second-order rate of user Ulk for FBL coding along with the normal approximation (NA) is given by
@@
where $n_{t}$ is the packet length, ${\bf S}_{lk}={\bf H}_{lk,l}{\bf P}_{lk}{\bf H}_{lk,l}^{H}$ is the covariance matrix of the desired signal at the user Ulk, while ${\bf D}_{lk}$ is the covariance matrix of the interfering signals plus noise, given by
@@
where ${\bf\Lambda}_{lk}=\text{diag}\left(\lambda_{lk1},\lambda_{lk2},\cdots,\lambda_{lkI}\right)$ is a diagonal matrix, containing the non-zero eigenvalues of the positive semidefinite (PSD) matrix ${\bf D}^{-1}_{lk}{\bf S}_{lk}$, and $I\leq\min(N_{BS,l},N_{u,lk})$ is equal to the rank of ${\bf H}_{lk,l}{\bf P}_{lk}{\bf H}_{lk,l}^{H}$, which also represents the number of parallel channels. The parameter $\lambda_{lki}$ is actually the signal-to-interference-plus-noise ratio (SINR) at the $i$-th parallel channel of user Ulk. Finally, $V_{lk}=\sum_{i=1}^{I}V_{lki}$ is the channel dispersion of Ulk, where $V_{lki}$ is the channel dispersion of the $i$-th parallel channel of user Ulk.
@@
Remark 1.
The reliability constraint is modeled by utilizing the maximum tolerable error rate, $\epsilon$, in the FBL rates. Moreover, the coding length $n_{t}$ should be proportional to the tolerable latency. Indeed, a more stringent latency constraint leads to employing a shorter block length [55, 56]. Additionally, the latency constraint can be modeled as a constraint on the minimum rate as discussed in [57, Remark 1], [15, Sec. II.D], [58].
The EE of Ulk is defined as [59]
@@
where $\eta^{-1}$ is the power efficiency of the transmit devices at the BSs, $P_{c}$ is the constant power consumption of the system (including the devices of the BSs, RISs and Ulk) to transmit data to a user, which is given by [28, Eq. (27)], and $r_{lk}$ is given by Lemma 1. Note that to compute $P_{c}$, the constant power of the devices of the BSs and RISs is normalized by the number of users served. Moreover, the global EE (GEE) of the network is defined as [59]
@@
which quantifies how energy efficient the network is. Finally, the transmission delay of Ulk upon transmitting a packet with length $n_{t}$ is $d_{lk}=\frac{n_{t}}{r_{lk}}$.
@@ II-D Problem Statement
where constraint (24c) can be interpreted as a latency constraint for each user, as discussed in [14, Sec. II.D]. Moreover, functions $f_{i}$, $\forall i$ are, in general, non-linear functions of the optimization variables. These functions can be, but are not restricted to, linear functions of the rates/EEs and/or transmit/receive powers. For instance, $f_{i}$ can be a function of the sum rate, minimum rate/EE, transmission/receive power, interference temperature at a user, transmission delay and so on. Additionally, $f_{i}$ can be a non-linear function of the rates/EEs such as the geometric mean of the rates as in [60]. Note that (24) may also include minimization problems such as the total power minimization subject to a given target rate, maximum transmission delay minimization, and interference temperature minimization. In this case, $f_{i}$ can be chosen as, e.g., $-\sum_{l}\text{Tr}({\bf P}_{l})$ or $-\max_{\forall lk}\{d_{lk}\}$. Therefore, the general problem in (24) can include an extensive range of optimization scenarios, encompassing the maximization of the minimum weighted rate, sum rate, global EE and minimum EE. We refer the reader to [36, Sec. II.B] for more discussions on the format of the functions $f_{i}$s as well as of the family of optimization problems that can be formulated as (24).
@@ III Proposed optimization framework
In this section, we propose iterative schemes for solving (24) by leveraging AO, MM-based, and FP algorithms. Specifically, we first fix the RIS coefficients to $\{{\bf\Psi}^{(t-1)}\}$ and update the transmit covariance matrices as $\{{\bf P}^{(t)}\}$ by solving (24). We then alternate and update the RIS coefficients, while $\{{\bf P}\}$ is fixed to $\{{\bf P}^{(t)}\}$. We iterate this procedure until convergence is reached. Unfortunately, the optimization problems are non-convex and complicated even when the RIS elements (or covariance matrices) are fixed. Thus, we propose a suboptimal scheme based on MM to solve the corresponding problems. Below, we present our solutions for updating the transmit covariance matrices and RIS elements in separate subsections.
@@ III-A Updating Transmit Covariance Matrices
To update $\{{\bf P}\}$, we introduce a new set of variables $\{{\bf Q}\}=\{{\bf Q}_{lk}\}_{\forall lk}$, where ${\bf Q}_{lk}$ is a positive semi-definite matrix and ${\bf P}_{lk}={\bf Q}_{lk}{\bf Q}_{lk}^{H}$. Equivalently, we can compute ${\bf Q}_{lk}$ as ${\bf P}_{lk}^{1/2}$. To attain a suboptimal solution for (24), we leverage an MM-based technique. More specifically, we first obtain suitable concave surrogate functions for the rates. Then, we update ${\bf P}_{lk}^{(t)}={\bf Q}_{lk}^{(t)}{\bf Q}_{lk}^{(t)^{H}}$, $\forall l,k$, by solving the corresponding surrogate optimization problems. To derive concave lower bounds for the FBL rates, we utilize the bounds in the following lemmas.
@@
$r_{lk}\geq\tilde{r}_{lk}=a_{lk}+2\sum_{ij}\mathfrak{R}\left\{\text{{\em Tr}}\left({\bf A}_{lk,ij}\mathbf{Q}_{ij}^{H}\bar{\mathbf{H}}_{lk,i}^{H}\right)\right\}\\ -\text{{\em Tr}}\left({\bf B}_{lk}(\mathbf{H}_{lk,l}\mathbf{Q}_{lk}\mathbf{Q}_{lk}^{H}\mathbf{H}_{lk,l}^{H}+\mathbf{D}_{lk})\right)$ (28)
@@
$a_{lk}$ $\!=\!\ln\left|\!{\bf I}\!+\!\bar{\bf D}_{lk}^{-1}\bar{\bf S}_{lk}\!\right|\!\!-\!\!\text{{\em Tr}}\!\left(\bar{\bf D}_{lk}^{-1}\bar{\bf S}_{lk}\right)\!\!-\!\frac{Q^{-1}\!(\epsilon)\!({\bar{V}_{lk}}\!\!+\!2I)}{2\sqrt{n_{t}\bar{V}_{lk}}},$
@@
${\bf B}_{lk}$ $\!=\bar{\mathbf{D}}^{-1}_{lk}\!-\!(\bar{\mathbf{S}}_{lk}+\bar{\mathbf{D}}_{lk})^{-1}\!\!$
$\hskip 17.07182pt+\!\frac{Q^{-1}(\epsilon)}{\sqrt{n_{t}\bar{V}_{lk}}}(\bar{\mathbf{S}}_{lk}+\bar{\mathbf{D}}_{lk})^{-1}\bar{\mathbf{D}}_{lk}(\bar{\mathbf{S}}_{lk}+\bar{\mathbf{D}}_{lk})^{-1}\!,$
@@
We denote the surrogate functions for $f_{i}$ by $\tilde{f}_{i}$, which are obtained by substituting the concave lower bounds $\tilde{r}_{lk}$ in (24). For instance, if ${f}_{i}$ is equal to the sum rate, then $\tilde{f}_{i}=\sum_{\forall lk}\tilde{r}_{lk}$. Moreover, if ${f}_{i}$ represents the EE of Ulk, then we have
@@
which is a fractional function of $\{{\bf Q}\}$ with a concave numerator and convex denominator. Moreover, if ${f}_{i}$ is a function of the transmission delay, then we have $\tilde{f}_{i}=-\frac{n_{t}}{\tilde{r}_{lk}}$. Note that although the surrogate lower bounds for the rates in Lemma 5 are concave, the $\tilde{f}_{i}$s are not necessarily concave, since they might be a linear function of the EE metrics. Substituting the $f_{i}$s by the $\tilde{f}_{i}$s leads to
@@
The optimization problem (33) is convex for the maximization of the minimum and/or sum rates. Hence, it can be efficiently solved by existing numerical tools. Note that our framework can also optimize other SE metrics, such as the geometric mean of users. As shown in [60], the maximization of the geometric mean of the users can be solved by solving a sequence of weighted sum rate maximization problems, which can be efficiently handled by our framework.
Unfortunately, (33) is non-convex for GEE maximization as well as for the maximization of the minimum weighted EE of the users, since the EE and/or GEE functions are not concave in $\{{\bf Q}\}$. Fortunately, a solution of the minimum weighted EE of the users and/or GEE maximization problems can be obtained by Dinkelbach-based algorithms, since the numerator of $\tilde{e}_{lk}$ (or $\tilde{g}$) is concave, while its denominator is convex. The problem in (33) is not convex, when minimizing the maximum transmission delay of the users, which can be considered as a latency metric from a physical layer point of view. In the following, we solve (33) for the maximization of the minimum weighted EE of the users, the maximization of the GEE and the minimization of the maximum transmission delay.
@@ III-A1 Maximization of the Minimum EE
In this case, (33) can be written as
@@
Upon employing the generalized Dinkelbach algorithm (GDA), we can derive the globally optimal solution of (34) by iteratively solving the convex optimization problem [59]
@@ III-A2 Maximization of the GEE
In this case, (33) is equivalent to
@@
III-A3 Minimization of the maximum transmission delay
To minimize the maximum delay, we have to solve
[…]
which is equivalent to maximizing the minimum rate of users as
[…]
which is a convex optimization problem. Note that the transmission delay $d_{lk}$ is a monotonically decreasing function of $r_{lk}$, and it is minimized when $r_{lk}$ is maximized, which makes the solution of (38) equivalent to (39).
III-A4 Discussion on Single-stream Data Transmission
In this case, the rank of matrix ${\bf P}_{lk}$ for all $l,k$ is equal to one. This means that matrix ${\bf P}_{lk}$ can be written as ${\bf P}_{lk}={\bf q}_{lk}{\bf q}_{lk}^{H}$, where ${\bf q}_{lk}\in\mathbb{C}^{N_{BS,l}\times 1}$. In other words, when single-stream data transmission is employed, the BSs perform only beamforming to transmit data, and we have to optimize the beamforming vectors instead of transmit covariance matrices. Therefore, the computational complexity of single-stream data transmission is lower than that of employing multiple streams. However, this lower computational complexity is attained at the cost of a significant performance loss, especially when the maximum number of streams, i.e., $\min(N_{BS,l},N_{u,lk})$ for Ulk, increases. Indeed, as the network size increases, more advanced transmission and resource allocation techniques, involving higher complexities, are needed to provide satisfactory performance.
Note that if the transmitter and/or the receiver are equipped with only a single antenna, we are restricted to single-stream data transmissions. Obviously, the single-stream scheme becomes increasingly suboptimal, when the maximum number of streams grows. To derive single-stream data transmission solutions, we only have to replace matrices ${\bf Q}_{lk}$ by vectors ${\bf q}_{lk}$ and employ the lower bounds in Lemma 5. Indeed, the schemes proposed in this subsection can be applied for both single- and multiple-stream data transmission.
@@ III-B Optimizing the RIS Elements
Now, we update $\{\bf{\Psi}\}$ by solving (24) for fixed $\{{\bf P}^{(t)}\}$, i.e.,
@@
which is non-convex since the rates are not concave in $\{\bf{\Psi}\}$ and the set $\mathcal{T}$ might be non-convex. Note that as discussed in Section III-A, the minimization of the maximum transmission delay is equivalent to the maximization of the minimum rate. Hence, our focus in this subsection is on the SE and EE metrics. In the following, we first consider reflective RISs and then, describe how the solution can be applied to the more sophisticated STAR-RIS, utilizing the MS scheme.
III-B1 Reflective RISs
To find a suboptimal solution for (40), we leverage an approach based on MM. Specifically, we first obtain a suitable concave lower bound for the rates and then, convexify $\mathcal{T}$ if it is not already a convex set. Since the rates have similar structures in $\{{\bf Q}\}$ and in $\{\bf{\Psi}\}$, we can utilize the concave lower bounds in Lemma 5 to attain concave lower bounds for the rates and construct suitable surrogate optimization problems for updating $\{\bf{\Psi}\}$. In the subsequent corollary, we provide the concave lower bounds for the rates.The proof of this corollary closely resembles that of Lemma 5, and thus, we omit it here.